x dy/dx = y - x


1
Expert's answer
2021-02-24T14:26:23-0500

Solution.

"x\frac{dy}{dx}=y-x."

This is linear inhomogeneous differential equation of the 1st order.


"xy'-y=-x,\newline
y'-\frac{1}{x}y=-1."

The solution of the equation is written in the form


"y=y_c+y_p,"

where "y_c" is the solution of homogeneous equation, and "y_p" is a particular solution of the inhomogeneous equation.

"y'-\frac{1}{x}y=0,"

"\frac{dy}{y}=\frac{dx}{x},"

"\int\frac{dy}{y}=\int\frac{dx}{x},"

"\ln{|y|}=\ln{|x|}+C_1,"

"|y|=|x|e^{C_1},"

from here "y_c=Cx," where "C" is some constant, "C\neq 0."

"y_p=u(x)x,"

"u'x+u-\frac{1}{x}ux=-1,"

"u'x=-1,"

"\frac{du}{dx}=-\frac{1}{x},"

"u=-\ln|x|," from here

"y_p=-\ln{|x|}x."

So, "y=Cx-x\ln{|x|}."

Answer. "y=Cx-x\ln{|x|}."